# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v1

@@ Abstract
We introduce a novel sampler called the energy based diffusion generator for generating samples from arbitrary target distributions. The sampling model employs a structure similar to a variational autoencoder, utilizing a decoder to transform latent variables from a simple distribution into random variables approximating the target distribution, and we design an encoder based on the diffusion model. Leveraging the powerful modeling capacity of the diffusion model for complex distributions, we can obtain an accurate variational estimate of the Kullback-Leibler divergence between the distributions of the generated samples and the target. Moreover, we propose a decoder based on generalized Hamiltonian dynamics to further enhance sampling performance. Through empirical evaluation, we demonstrate the effectiveness of our method across various complex distribution functions, showcasing its superiority compared to existing methods.
@@ 1 Introduction
In various fields such as computational chemistry, statistical physics, and machine learning, the challenge of sampling from a Boltzmann distribution corresponding to a high-dimensional and complex energy function is ubiquitous. Unlike training tasks for data-driven generative models, where existing data can be leveraged to learn complex distributions, sampling from such Boltzmann distributions poses a distinct and formidable challenge due to the absence of readily available data. This presents a complex and difficult problem that has yet to be effectively addressed.
@@
Markov Chain Monte Carlo (MCMC) methods have offered a pivotal solution to the challenge of sampling from high-dimensional distributions. These methods operate by iteratively generating candidates and updating samples, ultimately achieving asymptotic unbiasedness in the limit of infinite sampling steps (Hastings, 1970). To further amplify the sampling efficiency of MCMC, the integration of Brownian dynamics and Hamiltonian dynamics (Roberts & Tweedie, 1996; Neal et al., 2011; Betancourt et al., 2017; Cheng et al., 2018) has been introduced, providing a means to generate new samples. This integration results in improved acceptance probabilities and accelerated convergence rates. In recent years, researchers have proposed trainable transition kernels as a strategy for generating candidate samples, showcasing notable advancements in augmenting the efficiency and effectiveness of the sampling process (Song et al., 2017; Levy et al., 2017; Liu & Sun, 2022). However, the prolonged mixing time of MCMC still constrains its performance on high-dimensional complex distributions in practical applications.
@@
Variational inference (VI) is another crucial approach for addressing intractable distributions. VI leverages a generator capable of swiftly producing samples to approximate the target Boltzmann distribution, and the optimization of the generator’s parameters is then undertaken to minimize a statistical distance, such as the Kullback-Leibler (KL) divergence, between the distribution of the generated samples and the target distribution. Due to its ability to model complex distributions and provide explicit probability density functions, normalizing flow (NF) has been extensively applied to construct generators for VI methods (Li & Wang, 2018; Noé et al., 2019; Köhler et al., 2020; Boyda et al., 2021; Vaitl et al., 2022; Albergo et al., 2022; Gerdes et al., 2022; Felardos, 2022; van Leeuwen et al., 2023; Plainer et al., 2023). However, the bijective nature of NFs imposes a constraint on their effective capacity, often rendering it insufficient for certain sampling tasks. When the probability density function of the generated sample distribution is unknown, leveraging Stein Discrepancy provides a means to construct the gradient flow of KL divergence for refining sample positions (Liu & Wang, 2016; Liu, 2017). However, these methods often involve computationally expensive tasks, requiring the computation of kernel functions and their gradients. In addition, the combination of MCMC and VI methods stands as a current focal point in research (Salimans et al., 2015; Zhang & Hernández-Lobato, 2018; Habib & Barber, 2018; Ruiz & Titsias, 2019; Wu et al., 2020; Shen et al., 2021). This combination seeks to harness the strengths of both approaches, offering a promising avenue for addressing the challenges associated with sampling from high-dimensional distributions and enhancing the efficiency of probabilistic modeling.
@@
With the prosperity of diffusion based generative models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; 2021), they have been applied to address challenges in the sampling problem. By training time-dependent score matching neural networks, methods proposed in (Zhang & Chen, 2021; Berner et al., 2022; Vargas et al., 2023) shape the Gaussian distribution into complex target densities, employing the KL divergence as the loss function. To mitigate mode-seeking issues, (Richter et al., 2023) introduces the log-variance loss, showcasing favorable properties. Additionally, an alternative training objective is outlined in (Máté & Fleuret, 2023), relying on flexible interpolations of the energy function and demonstrating substantial improvements for multi-modal targets. However, a common drawback of these methods is their dependence on numerical differential equation solvers to compute time integrals, which can incur both significant computational costs and time constraints.
@@
In this research endeavor, we present a novel approach termed the energy based diffusion generator (EDG), drawing inspiration from both VI techniques and the diffusion model. The architecture of EDG closely resembles that of a variational autoencoder (VAE) (Kingma et al., 2019), comprising a decoder and an encoder. The decoder is flexible in mapping latent variables distributed according to tractable distributions to samples, without the imposition of constraints such as bijectivity, and we design in this work a decoder based on generalized Hamiltonian dynamics to enhance sampling efficiency. The encoder utilizes a diffusion process, enabling the application of score matching techniques for precise and efficient modeling of the conditional distribution of latent variables given samples. Unlike existing diffusion based methods, the loss function of EDG allows for the convenient computation of unbiased estimates in a mini-batch fashion, eliminating the need for numerical solutions to ordinary differential equations (ODEs) or stochastic differential equations (SDEs) during training. Numerical experiments conclusively demonstrate the effectiveness of EDG.
@@ 2 Preliminaries and setup
where the normalizing constant $Z=\int\exp(-U(x))\mathrm{d}x$ is usually computationally intractable. To tackle this challenge, the Boltzmann generator (Noé et al., 2019), along with its various extensions (Felardos, 2022; van Leeuwen et al., 2023; Plainer et al., 2023), has emerged as a prominent technique in recent years. These approaches harness NF to parameterize a trainable and analytical density function, with parameter optimization achieved through the minimization of the KL divergence between the surrogate density and $\pi$. Nevertheless, in contrast to typical generative models, the pursuit of exact probability density computation imposes substantial constraints on NFs: Each transformation layer must be a bijection, and the determinant of its Jacobian matrix can be tractably computed. These requirements inherently limit the capacity of NFs to effectively model complex distributions.
@@
Here, $z_{0}\sim p_{D}(z_{0})$ is a tractable latent variable with a known prior distribution, often a standard multivariate normal distribution, and $\phi$ denotes the parameter of the decoder. Similar to the VAE, we can formulate $p_{D}(x|z_{0};\phi)$ as $\mathcal{N}(x|\mu(z_{0};\phi),\Sigma(z_{0};\phi))$, where both $\mu$ and $\Sigma$ are represented by neural networks (NNs).
While the marginal distribution $p_{D}(x)$ of generator samples remains intractable, we can derive an upper bound for the KL divergence:
@@
$=$ $\mathbb{E}_{p_{D}(z_{0})\cdot p_{D}(x|z_{0},\phi)}\left[\log\frac{p_{D}(z_{0})p_{D}(x|z_{0},\phi)}{p_{E}(z_{0}|x,\theta)}+U(x)\right]+\log Z.$
@@
It seems that we have only increased the complexity of the problem, as we are still required to approximate a conditional distribution. However, in the upcoming section, we demonstrate that we can effectively construct the encoder using the diffusion model (Sohl-Dickstein et al., 2015; Song et al., 2020) and optimize all parameters without the need to numerically solve ordinary or stochastic differential equations.
3 Energy based diffusion generator
Motivated by the diffusion model (Song et al., 2020; 2021), we now discuss the idea of the energy based diffusion generator (EDG). In the EDG framework, we initiate a diffusion process from the latent variable $z_{0}$ and combine it with the decoder, resulting in what we term the “decoding process”:
$z_{0}\in R^{D}\sim p_{D}(z_{0}),\quad x|z_{0}\sim p_{D}(x|z_{0};\phi)$
@@ 3 Energy based diffusion generator
Now, let’s explore a scenario where the initial distribution of $z_{0}$ in (2) is defined as $p_{D}(z_{0}|x)\propto p_{D}(z_{0})\cdot p_{D}(x|z_{0})$. In this context, we can derive the following reverse-time diffusion equation (Anderson, 1982):
@@
where $\tilde{t}=T-t$ denotes the reverse time. Since the independence between $z_{T}$ and $z_{0}$ also implies the independence between $z_{T}$ and $x$, we can simulate (3) by initializing it from the marginal distribution of $p_{D}(z_{T})$ to acquire samples of $p_{D}(z_{0}|x)$. Nonetheless, practical implementation of this simulation presents challenges due to the intractability of the score function $\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}}|x)$. To address this challenge, we resort to approximating the score function using a parametric model denoted as $s(z_{\tilde{t}},x,\tilde{t};\theta)$. This approximation leads to what we refer to as the “encoding process”, achieved by integrating the parametric reverse-time diffusion process and the target distribution of $x$:
@@
$\mathrm{d}z_{\tilde{t}}=-\left(f(z_{\tilde{t}},\tilde{t})-g(\tilde{t})^{2}s(z_{\tilde{t}},x,\tilde{t}|\theta)\right)\mathrm{d}\tilde{t}+g(\tilde{t})\mathrm{d}W_{\tilde{t}},\quad\tilde{t}=T-t.$ (4)
@@
Fig. 1 offers a visual depiction of the decoding and encoding processes.
@@
3.1 Loss function
To optimize parameters of the decoding and encoding processes, we can minimize the KL divergence between the joint distributions of $(x,z_{[\cdot]})$ provided by the decoding and encoding processes, where $z_{[\cdot]}$ denotes the path $\left\{z_{t}\right\}_{t\in[0,T]}$ from 00 to $T$. The divergence is also an upper bound of $D_{KL}(p_{E}(x)||\pi(x))$ like that in (1) due to the data processing inequality, and can be expressed as
where
$\mathcal{L}(\theta,\phi)=\,$ $\mathbb{E}_{p_{D}}\left[\log p_{D}(x|z_{0};\phi)+U(x)\right]$
$+\int_{0}^{T}\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}\left[\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\nabla_{z_{t}}\cdot s\left(z_{t},x,t;\theta\right)+\left\|\nabla_{z_{t}}\log p_{D}(z_{t})\right\|^{2}\right]\mathrm{d}t.$ (6)
By utilizing the importance-sampled integration and Hutchinson estimator, we can obtain an equivalent expression of $\mathcal{L}(\theta,\phi)$ which can be efficiently and unbiasedly estimated by a small batch of samples of $p_{D}$:
where
$p(\epsilon)$ is the Rademacher distribution with $\mathbb{E}[\epsilon]=0,\mathrm{cov}(\epsilon)=I$, $p(t)$ is a proposal distribution of $t\in[0,T]$, and the weighting function $\lambda(t)=\frac{g(t)^{2}}{2p(t)}$.
The selection of $p(t)$ significantly influences the stochastic approximation of the second term on the r.h.s of (7). To mitigate the variance inherent in this approximation, we adopt a strategic approach outlined in (Choi et al., 2022), wherein we dynamically adjust the proposal distribution based on the historical buffer encompassing the $B$ most recent loss values concerning $t$ within the second term of (7).
For a comprehensive understanding of the loss function and the update strategy for $p(t)$, please refer to Appendix C, where detailed derivations are provided.
3.2 Boundary condition-guided encoder formulation
@@
$s(z,x,t;\theta)$ $=$ $\left(1-\frac{t}{T}\right)\cdot\nabla_{z_{0}}\left[\log p_{D}(x|z_{0}=z)+\log p_{D}(z_{0}=z)\right]+\frac{t}{T}\cdot\nabla_{z_{T}}\log p_{D}(z_{T}=z)$
$+\frac{t}{T}\left(1-\frac{t}{T}\right)s^{\prime}(z,x,t;\theta),$
@@
3.3 Computation of marginal encoder density
In line with the proof detailed in Appendix D of (Song et al., 2020), we can assert that the temporal evolution of $p_{E}(z_{t}|x)$ within the reverse-time diffusion process (4) adheres to the probability flow ordinary differential equation (ODE):

### v2

@@ Abstract
Sampling from Boltzmann distributions, particularly those tied to high-dimensional and complex energy functions, poses a significant challenge in many fields. In this work, we present the Energy-Based Diffusion Generator (EDG), a novel approach that integrates ideas from variational autoencoders and diffusion models. EDG leverages a decoder to transform latent variables from a simple distribution into samples approximating the target Boltzmann distribution, while the diffusion-based encoder provides an accurate estimate of the Kullback-Leibler divergence during training. Notably, EDG is simulation-free, eliminating the need to solve ordinary or stochastic differential equations during training. Furthermore, by removing constraints such as bijectivity in the decoder, EDG allows for flexible network design. Through empirical evaluation, we demonstrate the superior performance of EDG across a variety of complex distribution tasks, outperforming existing methods.
keywords:
Boltzmann distribution , Energy-based model , Generative model , Diffusion model , Variational autoencoder
[…]
@@ 1 Introduction
In various fields such as computational chemistry, statistical physics, and machine learning, the challenge of sampling from a Boltzmann distribution corresponding to a high-dimensional and complex energy function is ubiquitous [1, 2]. Unlike training tasks for data-driven generative models, where pre-sampled data can be utilized to learn complex distributions, sampling from Boltzmann distributions presents a unique and significant challenge due to the lack of readily available data [3, 4]. For example, simulating the phase transition of the Ising model can be framed as a sampling problem given the energy function, which presents a complex and difficult problem that has yet not to be effectively addressed [5, 6].
@@
Markov Chain Monte Carlo (MCMC) methods [7], along with Brownian and Hamiltonian dynamics [8, 9, 10, 11], have offered a pivotal solution to the challenge of sampling from high-dimensional distributions. These methods operate by iteratively generating candidates and updating samples, ultimately achieving asymptotic unbiasedness at the limit of infinite sampling steps. In recent years, researchers have proposed adaptive MCMC as a strategy for generating candidate samples, showcasing notable advancements in augmenting the efficiency and effectiveness of the sampling process [12, 13, 14]. However, the prolonged mixing time of MCMC still constrains its performance. Several studies have shown that using neural networks to construct and optimize proposal distributions in MCMC can significantly improve its efficiency [13, 15, 16]. However, there is still a lack of an effective and widely adaptable loss function to facilitate this optimization.
@@
Variational inference (VI) is another crucial approach for addressing intractable distributions. VI leverages a generator capable of swiftly producing samples to approximate the target Boltzmann distribution, and the optimization of the generator’s parameters is then undertaken to minimize a statistical distance, such as the Kullback-Leibler (KL) divergence, between the distribution of the generated samples and the target distribution. Due to its ability to model complex distributions and provide explicit probability density functions, normalizing flow (NF) has been extensively applied to construct generators for VI methods [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. However, the bijective nature of NFs imposes a constraint on their effective capacity, often rendering it insufficient for certain sampling tasks. Given the target density function and the generated samples, Stein Discrepancy [27, 28] provides a different means to evaluate the goodness of fit, while the computation of kernel functions and their gradients limits its performance in high-dimensional tasks. In addition, the combination of MCMC and VI methods stands as a current focal point in research [29, 30, 31, 32, 33, 34]. This combination seeks to harness the strengths of both approaches, offering a promising avenue for addressing the challenges associated with sampling from high-dimensional distributions and enhancing the efficiency of probabilistic modeling.
@@
With the prosperity of diffusion based generative models [35, 36, 37, 38], they have been applied to address challenges in the sampling problem. By training time-dependent score matching neural networks, methods proposed in [39, 40, 41] shape the Gaussian distribution into complex target densities, employing the KL divergence as the loss function. To mitigate mode-seeking issues, [42] introduces the log-variance loss, showcasing favorable properties. Additionally, an alternative training objective is outlined in [43], relying on flexible interpolations of the energy function and demonstrating substantial improvements for multi-modal targets. However, a common drawback of these methods is their reliance on numerical differential equation solvers for computing time integrals, which can lead to substantial computational costs.
@@
In this research endeavor, we present a novel approach termed the energy based diffusion generator (EDG), drawing inspiration from both the variational autoencoder (VAE) technique [44] and the diffusion model. The architecture of EDG closely resembles that of VAE, comprising a decoder and an encoder. The decoder is flexible in mapping latent variables distributed according to tractable distributions to samples, without the imposition of constraints such as bijectivity, and we design in this work a decoder based on generalized Hamiltonian dynamics to enhance sampling efficiency. The encoder utilizes a diffusion process, enabling the application of score matching techniques for precise and efficient modeling of the conditional distribution of latent variables given samples. Unlike existing diffusion-based methods, the loss function of EDG facilitates the convenient computation of unbiased estimates in a random mini-batch manner, removing the need for numerical solutions to ordinary differential equations (ODEs) or stochastic differential equations (SDEs) during training. Numerical experiments conclusively demonstrate the effectiveness of EDG.
@@ 2 Preliminaries and setup
where the normalizing constant $Z=\int\exp(-U(x))\mathrm{d}x$ is usually computationally intractable. To tackle this challenge, the Boltzmann generator [18], along with its various extensions [24, 25, 26], has emerged as a prominent technique in recent years. These approaches harness NF to parameterize a trainable and analytical density function, with parameter optimization achieved through the minimization of the KL divergence between the surrogate density and $\pi$. Nevertheless, in contrast to typical generative models, the pursuit of exact probability density computation imposes substantial constraints on NFs: Each transformation layer must be a bijection, and the determinant of its Jacobian matrix can be tractably computed. These requirements inherently limit the capacity of NFs to effectively model complex distributions.
@@
where $z_{0}\sim p_{D}(z_{0})$ is a latent variable drawn from a known prior distribution, typically a standard multivariate normal distribution. The parameter $\phi$ characterizes the decoder, and we define $p_{D}(x|z_{0};\phi)$ as a Gaussian distribution $\mathcal{N}(x|\mu(z_{0};\phi),\Sigma(z_{0};\phi))$, with both $\mu$ and $\Sigma$ parameterized by neural networks (NNs). Similar to the VAE, we aim to train the networks $\mu$ and $\Sigma$ such that the marginal distribution $p_{D}(x)$ of the generated samples aligns with the target distribution.
It is important to note that, unlike conventional data-driven VAEs, we do not have access to samples from the target distribution $\pi(x)$. In fact, obtaining such samples is precisely the goal of the generator. As a result, the variational approximation of the KL divergence $D_{KL}\left(\pi(x)||p_{D}(x)\right)$ cannot be used to train the model. Instead, in this work, we consider the following divergence and its upper bound:
@@
$=$ $\mathbb{E}_{p_{D}(z_{0})\cdot p_{D}(x|z_{0},\phi)}\left[\log\frac{p_{D}(z_{0})p_{D}(x|z_{0},\phi)}{p_{E}(z_{0}|x,\theta)}+U(x)\right]$
$+\log Z.$
@@
It seems that we have only increased the complexity of the problem, as we are still required to approximate a conditional distribution. However, in the upcoming section, we demonstrate that we can effectively construct the encoder using the diffusion model [35, 37] and optimize all parameters without the need to numerically solve ordinary or stochastic differential equations.
3 Energy-based diffusion generator
The diffusion model [37, 38] has emerged in recent years as a highly effective approach for estimating data distributions. Its core idea is to construct a diffusion process that progressively transforms data into simple white noise, and learn the reverse process to recover the data distribution from noise. In this work, we apply the principles of the diffusion model by incorporating a diffusion process into the latent space, enabling us to efficiently overcome the challenges arising from the variational framework for the sampling problem defined by Equation (1). We refer to the model produced by this method as the energy-based diffusion generator (EDG).
3.1 Model architecture
In the EDG framework, we initiate a diffusion process from the latent variable $z_{0}$ and combine it with the decoder, resulting in what we term the “decoding process”:
$z_{0}\in R^{D}\sim p_{D}(z_{0})\triangleq\mathcal{N}(x|0,I),\quad x|z_{0}\sim p_{D}(x|z_{0};\phi)$
@@ 3.1 Model architecture
If we only consider the statistical properties of the latent diffusion process $z_{[\cdot]}=\{z_{t}\}_{t\in[0,T]}$, it is uninformative and only describes a transition from one simple noise to another. However, when we account for the conditional distribution of $z_{t}$ given a sample $x$, the process $z_{[:]}$ represents the gradual transformation of the complex conditional distribution $p_{D}(z_{0}|x)\propto p_{D}(z_{0})\cdot p_{D}(x|z_{0})$ into the tractable distribution $p_{D}(z_{T}|x)=p_{D}(z_{T})$, where the independence between $z_{T}$ and $x$ results from the independence between $z_{0}$ and $z_{T}$ (see A). This implies that, starting from $z_{T}\sim p_{D}(z_{T})$, we can obtain samples from $p_{D}(z_{0}|x)$ by simulating the following reverse-time diffusion equation [45]:
@@
where $\tilde{t}=T-t$ denotes the reverse time. As in conventional diffusion models, practical implementation of this simulation is challenging due to the intractability of the score function $\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}}|x)$, and we therefore also use a neural network to approximate the score function, denoted as $s(z_{\tilde{t}},x,\tilde{t};\theta)$. This approximation leads to what we refer to as the “encoding process”, achieved by integrating the parametric reverse-time diffusion process and the target distribution of $x$:
@@
$\mathrm{d}z_{\tilde{t}}=-\left(f(z_{\tilde{t}},\tilde{t})-g(\tilde{t})^{2}s(z_{\tilde{t}},x,\tilde{t};\theta)\right)\mathrm{d}\tilde{t}+g(\tilde{t})\mathrm{d}W_{\tilde{t}},\quad\tilde{t}=T-t.$ (4)
@@
Fig. 1 offers a visual depiction of the decoding and encoding processes. It is important to highlight that latent diffusion models, which are primarily designed to address generative modeling in data-driven scenarios, which have recently gained significant attention [46, 47, 48]. Their main idea to use a pre-trained encoder and decoder to obtain a latent space that both effectively represents the data and facilitates efficient sampling, with the diffusion model learning the distribution of latent variables. Our model, EDG, leverages a similar idea to address energy-based sampling problems. The key differences between EDG and previous latent diffusion models in structure and algorithm are as follows: first, in EDG, the diffusion model itself functions as the encoder, eliminating the need for a separate encoder; second, by using a unified loss function, the decoder is trained jointly with the diffusion model (see Sec. 3.2).
@@
Below, we present the construction details of modules in EDG, which are used in our experiments. In practical applications, more effective neural networks can be designed as needed.
3.1.1 Boundary condition-guided score function model
@@
$s(z,x,t;\theta)$ $=$ $\left(1-\frac{t}{T}\right)\cdot\nabla_{z_{0}}\left[\log p_{D}(x|z_{0}=z)+\log p_{D}(z_{0}=z)\right]$
$+\frac{t}{T}\cdot\nabla_{z_{T}}\log p_{D}(z_{T}=z)+\frac{t}{T}\left(1-\frac{t}{T}\right)s^{\prime}(z,x,t;\theta),$
@@
3.1.2 Generalized Hamiltonian dynamics-based decoder
Inspired by generalized Hamiltonian dynamics (GHD) [12, 13], the decoder generates the output $x$ using the following process. First, an initial sample and velocity ($y,v$) are generated according to the latent variable $z_{0}$. Then, $(y,v)$ is iteratively updated as follows:
$v$ ::=:= $v-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right),$
$y$ ::=:= $y+\epsilon(l;\phi)\left(v_{k}\odot e^{\epsilon_{0}Q_{y}(v_{k},l;\phi)}+T_{y}(v_{k},l;\phi)\right),$
$v$ ::=:= $v-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right).$
Finally, the decoder output $x$ is given by:
$x=y-\exp(\epsilon_{0}\eta(y;\phi))\nabla U(y)+2\exp(\epsilon_{0}\eta(y;\phi))\xi$
where $\xi$ is distributed according to the standard Gaussian distribution. The equation can be interpreted as a finite-step approximation of Brownian dynamics $\mathrm{d}y=-\nabla U(y)\,\mathrm{d}t+\mathrm{d}W_{t}$. In the above procedure, $Q_{v}$, $T_{v}$, $Q_{y}$, $T_{y}$, $\epsilon$, and $\eta$ are all trainable neural networks.
The key advantages of the GHD-based decoder are twofold. First, it effectively leverages the gradient information of the energy function, and our experiments show that it can enhance the sampling performance for multi-modal distributions. Second, by incorporating trainable correction terms and steps into the classical Hamiltonian dynamics, it achieves a good decoder density with only a few iterations. The complete decoding process can be found in C.
3.2 Loss function
To optimize parameters of the decoding and encoding processes, we can minimize the KL divergence between the joint distributions of $(x,z_{[\cdot]})$ provided by the decoding and encoding processes. The divergence is also an upper bound of $D_{KL}(p_{D}(x)||\pi(x))$ like that in (1) due to the data processing inequality. As derived in B, the KL divergence can be expressed as
where